Company Belera is a manufacturing firm producing specialized equipments. June 1, 2018 B/S shows below balances:

Item Beg

RM 4,270

MOH Applied 32,740

MOH Incurred 36,650

June 1, 2018 subsidiary ledgers (job cost sheets) show 3 jobs are being performed with following balances :

Assigned Costs DM DL MOH

R 660 320 408

S 800 548 675

L 980 580 765

During June, RM purchases on account were TL4,940, and all wages (inclusive of payroll taxes) were paid in cash. Additional MOH of depreciation on factory equipment of TL900 was incurred together with miscellaneous production costs of TL440 on account. A summary of MRSs and LTTs shows below information.

Assigned Costs DM DL

R 870 850

K 2,000 870

S 560 360

L 1,300 1,260

R 360 390

General Use 1,500 1,208

MOH was charged to jobs at a rate of 1.85 per TL of DL costs. Jobs R, S and L were completed in June and sold for a total of TL19,400 in cash.

a) Compute the balance of WIP as at June 30.

b) Compute the balance of FG as at June 30.

c) Calculate the balance of MOH account as at June 30.
Answer #1
R K S L Total
Beginning Balance 1388.00 2023.00 2325.00 5736.00
Costs added
DM 1230.00 2000.00 560.00 1300.00 5090.00
DL 1240.00 870.00 360.00 1260.00 3730.00
MOH 2294.00 1609.50 666.00 2331.00 6900.50
Total Costs added 4764.00 4479.50 1586.00 4891.00 15720.50
Total Costs 6152.00 4479.50 3609.00 7216.00 21456.50

a) Balance of WIP = 4479.50 i.e. for Job K

b) Balance of FG = 0, since all jobs completed are sold

c) Actual Manufacturing Overhead = 36650+900+440+1500+1208 = $40698
Overhead applied = 32740+6900.50 = 39640.50
Balance of MOH = 40698-39640.5 = 1057.50 i.e. under applied overhead